# ChatBench: From Static Benchmarks to Human-AI Evaluation

### v1

@@ A.1 Pilots and Incentivization
While this bonus could help to improve ecological validity, there was a risk that the incentives result in users cheating on the study, such as by searching for the question on Google or ChatGPT. To mitigate this risk, first we repeatedly required users to acknowledge that they would not use external tools (Figures A3 and A7) and we said, “Compensation could be affected if we detect that you are using external tool.” Second, we ran a second medium-sized pilot with incentives, with 60 participants on the three math datasets, and we compared the results between Pilots 1 and 2 to see if Pilot 2 had unrealistic increases in accuracy that could not be explained by slightly more user effort.
@@ A.2 ChatBench
To protect the privacy of our participants, first we mapped each person’s Prolific ID to a new, randomly generated string of 10 letters and digits, checking that there were no collisions between individuals. The worker_id field in ChatBench contains these new strings, instead of their original worker ID on Prolific. Second, we checked for personally identifying information in the user-AI conversations. We used the European Union GDPR’s definition of personal data: [^5] https://gdpr.eu/eu-gdpr-personal-data/.
@@
For our main statistical analyses (Section 4), we only used data from the full study, and not from the two pilots. We furthermore filtered the data following the criteria we described in our preregistration, [^6] https://aspredicted.org/n84n-sn3f.pdf. such as only keeping the workers’ answers from their first assignment if they had multiple. While we could control from the Prolific interface that workers could not participate in our task multiple times, once they opened our app, they could start the study then be taken back to the beginning of the study flow (Figure 1) if they refreshed the app. If they did so, we would not want to keep their data after they refreshed, since their behavior on the second time around could be affected by what they already saw the first time. When a worker opened the app or refreshed, they received a new assignment, defined by their combination of model (one of 2), user-AI condition (one of 2), subject (one of 3), and question batch (one of 7 or 18). The probability that they would receive the same assignment twice if they refreshed was very low (less than 1%), so we could check for multiple assignments to test whether they refreshed, and we used timestamps to determine their first assignment. Ultimately, we found very few workers (3%) with multiple assignments, and for those workers, we kept their data from their first assignment.
@@
A.3 Task Interface
We provide screenshots of all of the pages in our user study interface, including the Introduction Page (Figure A3), Phase 1 Tutorial (Figure A4), Confidence Page (Figure A5), User-Alone Page (Figure A6), Phase 2 Instructions (Figure A7), Phase 2 Tutorial (Figure A8), User-AI Page (Figure 2), and Feedback Page (Figure A9).
{adjustwidth}
-1cm-1cm Dataset Model Comparison $\mathrm{Acc}_{1}$ $\mathrm{SE}_{1}$ $\mathrm{Acc}_{2}$ $\mathrm{SE}_{2}$ $z$-value $p$-value Elementary Math GPT-4o AI letter zero shot vs. UserAI direct to ai 0.73 0.04 0.92 0.02 -3.92 $<$0.01 Elementary Math GPT-4o AI letter zero shot vs. UserAI answer first 0.73 0.04 0.90 0.02 -3.43 $<$0.01 Elementary Math GPT-4o AI letter few shot vs. UserAI direct to ai 0.74 0.04 0.92 0.02 -3.83 $<$0.01 Elementary Math GPT-4o AI letter few shot vs. UserAI answer first 0.74 0.04 0.90 0.02 -3.34 $<$0.01 Elementary Math GPT-4o AI free text vs. UserAI direct to ai 0.99 0.01 0.92 0.02 3.03 $<$0.01 Elementary Math GPT-4o AI free text vs. UserAI answer first 0.99 0.01 0.90 0.02 4.04 $<$0.01 Elementary Math GPT-4o User alone vs. UserAI direct to ai 0.78 0.03 0.92 0.02 -4.21 $<$0.01 Elementary Math GPT-4o User alone vs. UserAI answer first 0.78 0.03 0.90 0.02 -3.52 $<$0.01 High School Math GPT-4o AI letter zero shot vs. UserAI direct to ai 0.51 0.05 0.70 0.04 -3.20 $<$0.01 High School Math GPT-4o AI letter zero shot vs. UserAI answer first 0.51 0.05 0.73 0.03 -3.92 $<$0.01 High School Math GPT-4o AI letter few shot vs. UserAI direct to ai 0.49 0.04 0.70 0.04 -3.57 $<$0.01 High School Math GPT-4o AI letter few shot vs. UserAI answer first 0.49 0.04 0.73 0.03 -4.33 $<$0.01 High School Math GPT-4o AI free text vs. UserAI direct to ai 0.85 0.03 0.70 0.04 3.14 $<$0.01 High School Math GPT-4o AI free text vs. UserAI answer first 0.85 0.03 0.73 0.03 2.73 $<$0.01 High School Math GPT-4o User alone vs. UserAI direct to ai 0.41 0.03 0.70 0.04 -5.88 $<$0.01 High School Math GPT-4o User alone vs. UserAI answer first 0.41 0.03 0.73 0.03 -7.03 $<$0.01 College Math GPT-4o AI letter zero shot vs. UserAI direct to ai 0.45 0.07 0.52 0.08 -0.61 0.54 College Math GPT-4o AI letter zero shot vs. UserAI answer first 0.45 0.07 0.52 0.07 -0.72 0.47 College Math GPT-4o AI letter few shot vs. UserAI direct to ai 0.44 0.07 0.52 0.08 -0.72 0.47 College Math GPT-4o AI letter few shot vs. UserAI answer first 0.44 0.07 0.52 0.07 -0.85 0.40 College Math GPT-4o AI free text vs. UserAI direct to ai 0.73 0.06 0.52 0.08 2.23 0.03 College Math GPT-4o AI free text vs. UserAI answer first 0.73 0.06 0.52 0.07 2.40 0.02 College Math GPT-4o User alone vs. UserAI direct to ai 0.28 0.04 0.52 0.08 -2.67 $<$0.01 College Math GPT-4o User alone vs. UserAI answer first 0.28 0.04 0.52 0.07 -3.10 $<$0.01 Conceptual Physics GPT-4o AI letter zero shot vs. UserAI direct to ai 0.91 0.03 0.84 0.03 1.74 0.08 Conceptual Physics GPT-4o AI letter zero shot vs. UserAI answer first 0.91 0.03 0.84 0.03 1.70 0.09 Conceptual Physics GPT-4o AI letter few shot vs. UserAI direct to ai 0.96 0.02 0.84 0.03 3.22 $<$0.01 Conceptual Physics GPT-4o AI letter few shot vs. UserAI answer first 0.96 0.02 0.84 0.03 3.22 $<$0.01 Conceptual Physics GPT-4o AI free text vs. UserAI direct to ai 0.97 0.02 0.84 0.03 3.62 $<$0.01 Conceptual Physics GPT-4o AI free text vs. UserAI answer first 0.97 0.02 0.84 0.03 3.63 $<$0.01 Conceptual Physics GPT-4o User alone vs. UserAI direct to ai 0.55 0.03 0.84 0.03 -6.48 $<$0.01 Conceptual Physics GPT-4o User alone vs. UserAI answer first 0.55 0.03 0.84 0.03 -6.69 $<$0.01 Moral Scenarios GPT-4o AI letter zero shot vs. UserAI direct to ai 0.71 0.05 0.79 0.03 -1.47 0.14 Moral Scenarios GPT-4o AI letter zero shot vs. UserAI answer first 0.71 0.05 0.78 0.04 -1.13 0.26 Moral Scenarios GPT-4o AI letter few shot vs. UserAI direct to ai 0.80 0.04 0.79 0.03 0.27 0.79 Moral Scenarios GPT-4o AI letter few shot vs. UserAI answer first 0.80 0.04 0.78 0.04 0.49 0.63 Moral Scenarios GPT-4o AI free text vs. UserAI direct to ai 0.72 0.05 0.79 0.03 -1.26 0.21 Moral Scenarios GPT-4o AI free text vs. UserAI answer first 0.72 0.05 0.78 0.04 -0.93 0.35 Moral Scenarios GPT-4o User alone vs. UserAI direct to ai 0.73 0.03 0.79 0.03 -1.54 0.12 Moral Scenarios GPT-4o User alone vs. UserAI answer first 0.73 0.03 0.78 0.04 -1.05 0.29
{adjustwidth}
-1cm-1cm Dataset Model Comparison $\mathrm{Acc}_{1}$ $\mathrm{SE}_{1}$ $\mathrm{Acc}_{2}$ $\mathrm{SE}_{2}$ $z$-value $p$-value Elementary Math Llama-3.1-8b AI letter zero shot vs. UserAI direct to ai 0.45 0.04 0.86 0.03 -8.58 $<$0.01 Elementary Math Llama-3.1-8b AI letter zero shot vs. UserAI answer first 0.45 0.04 0.90 0.02 -10.50 $<$0.01 Elementary Math Llama-3.1-8b AI letter few shot vs. UserAI direct to ai 0.43 0.03 0.86 0.03 -9.39 $<$0.01 Elementary Math Llama-3.1-8b AI letter few shot vs. UserAI answer first 0.43 0.03 0.90 0.02 -11.53 $<$0.01 Elementary Math Llama-3.1-8b AI free text vs. UserAI direct to ai 0.88 0.03 0.86 0.03 0.56 0.58 Elementary Math Llama-3.1-8b AI free text vs. UserAI answer first 0.88 0.03 0.90 0.02 -0.65 0.51 Elementary Math Llama-3.1-8b User alone vs. UserAI direct to ai 0.81 0.03 0.86 0.03 -1.26 0.21 Elementary Math Llama-3.1-8b User alone vs. UserAI answer first 0.81 0.03 0.90 0.02 -2.70 $<$0.01 High School Math Llama-3.1-8b AI letter zero shot vs. UserAI direct to ai 0.32 0.03 0.62 0.04 -6.14 $<$0.01 High School Math Llama-3.1-8b AI letter zero shot vs. UserAI answer first 0.32 0.03 0.64 0.04 -6.89 $<$0.01 High School Math Llama-3.1-8b AI letter few shot vs. UserAI direct to ai 0.30 0.02 0.62 0.04 -7.09 $<$0.01 High School Math Llama-3.1-8b AI letter few shot vs. UserAI answer first 0.30 0.02 0.64 0.04 -7.98 $<$0.01 High School Math Llama-3.1-8b AI free text vs. UserAI direct to ai 0.64 0.04 0.62 0.04 0.24 0.81 High School Math Llama-3.1-8b AI free text vs. UserAI answer first 0.64 0.04 0.64 0.04 -0.16 0.87 High School Math Llama-3.1-8b User alone vs. UserAI direct to ai 0.45 0.03 0.62 0.04 -3.37 $<$0.01 High School Math Llama-3.1-8b User alone vs. UserAI answer first 0.45 0.03 0.64 0.04 -3.93 $<$0.01 College Math Llama-3.1-8b AI letter zero shot vs. UserAI direct to ai 0.35 0.04 0.46 0.07 -1.37 0.17 College Math Llama-3.1-8b AI letter zero shot vs. UserAI answer first 0.35 0.04 0.48 0.07 -1.56 0.12 College Math Llama-3.1-8b AI letter few shot vs. UserAI direct to ai 0.30 0.04 0.46 0.07 -1.97 0.05 College Math Llama-3.1-8b AI letter few shot vs. UserAI answer first 0.30 0.04 0.48 0.07 -2.18 0.03 College Math Llama-3.1-8b AI free text vs. UserAI direct to ai 0.41 0.05 0.46 0.07 -0.57 0.57 College Math Llama-3.1-8b AI free text vs. UserAI answer first 0.41 0.05 0.48 0.07 -0.74 0.46 College Math Llama-3.1-8b User alone vs. UserAI direct to ai 0.40 0.04 0.46 0.07 -0.75 0.46 College Math Llama-3.1-8b User alone vs. UserAI answer first 0.40 0.04 0.48 0.07 -0.93 0.35 Conceptual Physics Llama-3.1-8b AI letter zero shot vs. UserAI direct to ai 0.53 0.05 0.67 0.04 -2.25 0.02 Conceptual Physics Llama-3.1-8b AI letter zero shot vs. UserAI answer first 0.53 0.05 0.73 0.04 -3.22 $<$0.01 Conceptual Physics Llama-3.1-8b AI letter few shot vs. UserAI direct to ai 0.57 0.04 0.67 0.04 -1.64 0.10 Conceptual Physics Llama-3.1-8b AI letter few shot vs. UserAI answer first 0.57 0.04 0.73 0.04 -2.70 $<$0.01 Conceptual Physics Llama-3.1-8b AI free text vs. UserAI direct to ai 0.62 0.04 0.67 0.04 -0.77 0.44 Conceptual Physics Llama-3.1-8b AI free text vs. UserAI answer first 0.62 0.04 0.73 0.04 -1.80 0.07 Conceptual Physics Llama-3.1-8b User alone vs. UserAI direct to ai 0.46 0.03 0.67 0.04 -3.91 $<$0.01 Conceptual Physics Llama-3.1-8b User alone vs. UserAI answer first 0.46 0.03 0.73 0.04 -4.97 $<$0.01 Moral Scenarios Llama-3.1-8b AI letter zero shot vs. UserAI direct to ai 0.40 0.03 0.72 0.04 -6.01 $<$0.01 Moral Scenarios Llama-3.1-8b AI letter zero shot vs. UserAI answer first 0.40 0.03 0.74 0.04 -7.42 $<$0.01 Moral Scenarios Llama-3.1-8b AI letter few shot vs. UserAI direct to ai 0.31 0.03 0.72 0.04 -7.35 $<$0.01 Moral Scenarios Llama-3.1-8b AI letter few shot vs. UserAI answer first 0.31 0.03 0.74 0.04 -8.86 $<$0.01 Moral Scenarios Llama-3.1-8b AI free text vs. UserAI direct to ai 0.49 0.03 0.72 0.04 -4.07 $<$0.01 Moral Scenarios Llama-3.1-8b AI free text vs. UserAI answer first 0.49 0.03 0.74 0.04 -5.15 $<$0.01 Moral Scenarios Llama-3.1-8b User alone vs. UserAI direct to ai 0.79 0.03 0.72 0.04 1.34 0.18 Moral Scenarios Llama-3.1-8b User alone vs. UserAI answer first 0.79 0.03 0.74 0.04 1.00 0.32
{adjustwidth}
-1cm-1cm Count Model Dataset Condition Answer Type GPT-4o College Math answer-first userAIAnswer 134 userAnswer 283 direct-to-AI userAIAnswer 116 userAnswer 121 Conceptual Physics answer-first userAIAnswer 317 userAnswer 425 direct-to-AI userAIAnswer 351 userAnswer 117 Elementary Math answer-first userAIAnswer 542 userAnswer 697 direct-to-AI userAIAnswer 462 userAnswer 122 High School Math answer-first userAIAnswer 539 userAnswer 689 direct-to-AI userAIAnswer 463 userAnswer 122 Moral Scenarios answer-first userAIAnswer 242 userAnswer 331 direct-to-AI userAIAnswer 398 userAnswer 135 Llama-3.1-8b College Math answer-first userAIAnswer 119 userAnswer 251 direct-to-AI userAIAnswer 115 userAnswer 123 Conceptual Physics answer-first userAIAnswer 315 userAnswer 428 direct-to-AI userAIAnswer 333 userAnswer 112 Elementary Math answer-first userAIAnswer 485 userAnswer 620 direct-to-AI userAIAnswer 462 userAnswer 123 High School Math answer-first userAIAnswer 477 userAnswer 610 direct-to-AI userAIAnswer 464 userAnswer 125 Moral Scenarios answer-first userAIAnswer 349 userAnswer 471 direct-to-AI userAIAnswer 229 userAnswer 81
@@ B.3 Additional analyses from user study
The following conversation occurred between a user called "You" and an AI called "Bot", in which "You" tried to use "Bot" to answer a question. Extract the following information about this conversation.
1. Is the first substantive user prompt a: question, statement, or other?
2. Is the first user question: a near-"exact" rephrasing of the question, a component of the question, an erroneous or misinterpreted aspect of the question, or other?
3. Is the first AI answer: correct, incorrect, or not yet providing an answer?
4. Is the last AI answer: correct, incorrect, or not yet providing an answer?
5. Are there more than one AI answer?
5. Anywhere in the course of the conversation, does AI correct the user?
6. Anywhere in the course of the conversation, does the user correct the AI by selecting a different answer than what the AI recommended?
7. Anything unusual or interesting about this interaction that you noticed?
@@
Here is what the user selected as their answer:
{selectedAnswer}
@@
We used the structured data from this analysis for a number of results, described in Section 4.3, such as how often user-AI interactions “mirror” AI benchmarks and how often AI provided the correct answer in the user-AI conversations. We also used this structured data to measure how often the user corrects the AI model’s mistake, by computing the fraction of user-AI interactions where the last AI answer in the conversation is wrong but the user still answered correctly (Figure B1). We find that users are much likelier to correct Llama-3.1-8b than GPT-4o, which helps to explain how some of the gap in the model’s AI-alone performance is closed in the user-AI setting.

### v2

@@ A.2 Pilots and Incentivization
While this bonus could help to improve ecological validity, there was a risk that the incentives result in users cheating on the study, such as by searching for the question on Google or ChatGPT. To mitigate this risk, first we repeatedly required users to acknowledge that they would not use external tools (Figures A4 and A8) and we said, “Compensation could be affected if we detect that you are using external tool.” Second, we ran a second medium-sized pilot with incentives, with 60 participants on the three math datasets, and we compared the results between Pilots 1 and 2 to see if Pilot 2 had unrealistic increases in accuracy that could not be explained by slightly more user effort.
@@ A.3 ChatBench
To protect the privacy of our participants, first we mapped each person’s Prolific ID to a new, randomly generated string of 10 letters and digits, checking that there were no collisions between individuals. The worker_id field in ChatBench contains these new strings, instead of their original worker ID on Prolific. Second, we checked for personally identifying information in the user-AI conversations. We used the EU GDPR’s definition of personal data: [^5] https://gdpr.eu/eu-gdpr-personal-data/.
@@
For our main statistical analyses (Section 4), we only used data from the full study, and not from the two pilots. We furthermore filtered the data following the criteria we described in our preregistration, [^6] https://aspredicted.org/n84n-sn3f.pdf. such as only keeping the workers’ answers from their first assignment if they had multiple. While we could control from Prolific that workers could not participate in our task multiple times, once they opened our app, they could start the study then be taken back to the beginning of the study flow (Figure 1) if they refreshed the app. If they did so, we would not want to keep their data after they refreshed, since their behavior on the second time around could be affected by what they already saw the first time. When a worker opened the app or refreshed, they received a new assignment, defined by their combination of model (one of 2), user-AI condition (one of 2), subject (one of 3), and question batch (one of 7 or 18). The probability that they would receive the same assignment twice if they refreshed was very low (less than 1%), so we could check for multiple assignments to test whether they refreshed, and we used timestamps to determine their first assignment. Ultimately, we found very few workers (3%) with multiple assignments, and for those workers, we kept their data from their first assignment.
@@
Count
Model Dataset Condition Answer Type
GPT-4o College Math answer-first userAIAnswer 134
userAnswer 283
direct-to-AI userAIAnswer 116
userAnswer 121
Conceptual Physics answer-first userAIAnswer 317
userAnswer 425
direct-to-AI userAIAnswer 351
userAnswer 117
Elementary Math answer-first userAIAnswer 542
userAnswer 697
direct-to-AI userAIAnswer 462
userAnswer 122
High School Math answer-first userAIAnswer 539
userAnswer 689
direct-to-AI userAIAnswer 463
userAnswer 122
Moral Scenarios answer-first userAIAnswer 242
userAnswer 331
direct-to-AI userAIAnswer 398
userAnswer 135
Llama-3.1-8b College Math answer-first userAIAnswer 119
userAnswer 251
direct-to-AI userAIAnswer 115
userAnswer 123
Conceptual Physics answer-first userAIAnswer 315
userAnswer 428
direct-to-AI userAIAnswer 333
userAnswer 112
Elementary Math answer-first userAIAnswer 485
userAnswer 620
direct-to-AI userAIAnswer 462
userAnswer 123
High School Math answer-first userAIAnswer 477
userAnswer 610
direct-to-AI userAIAnswer 464
userAnswer 125
Moral Scenarios answer-first userAIAnswer 349
userAnswer 471
direct-to-AI userAIAnswer 229
userAnswer 81
@@ B.3 Additional analyses from user study
The following conversation occurred between a user called "You" and an AI called "Bot", in which the user "You" was given a question and tried to use "Bot" to answer it correctly. Answer the following queries about this conversation.
1. True or false: the first user question is an exact or near-exact rephrasing of the question that the user was given, to which an intelligent person or AI would respond with the correct answer among the answer choices (without seeing anything else, including the answer choices if the user didnt supply them). Answer with "true" or "false".
2. True or false: the first time the AI answers the question, it is with a correct answer to the question that the user was given. Answer with "true" or "false".
3. True or false: the AI answers the question that the user was given more than once during the conversation. Answer with "true" or "false".
4: The first time the AI answers the question, which answer choice does its answer correspond to? Answer with "A", "B", "C", "D", or "none".
@@
Format your answer as a JSON object with the following keys: "1:first_user_question","2:first_ai_answer_correct","3:more_than_one_ai_answer","4:first_ai_answer_option".
@@
We used the structured data from this analysis for a number of results, described in Section 4.3, such as how often user-AI interactions “mirror” AI benchmarks and how often AI provided the correct answer in the user-AI conversations. We also used this structured data to measure how often the user corrects the AI model’s mistake, by computing the fraction of user-AI interactions where the first AI answer in the conversation is wrong but the user still answered correctly (Figure B1). We find that this occurs in a small fraction of cases, and users are generally more likely to correct Llama-3.1-8b than GPT-4o.
In Table B1, we provide additional statistics of the user-AI conversations, some computed from the structured data described above. Overall, we find that the conversations tend to be short in length, but each message is long and there is genuine information being exchanged. This is natural in a QA setting where we are only asking the user to answer one single question (no follow-ups) with the help of AI. Even without those constraints, user-AI conversations in the real world are also often short: for example, in WildChat Zhao et al. (2024), which includes 1 million user-AI conversations “in the wild”, around 60% of their conversations consist of only one user-AI turn.
In ChatBench, we also find that the vast majority of users are putting in effort to try to use AI to answer the question, as opposed to not putting in effort (e.g., “hello”, without a second message) or not using the AI (e.g., “i solved this already”). Oftentimes, this effort takes the form of the user providing a near-exact rephrasing of the question or a component of the question. The user’s effort is substantiated by what follows, since in the majority of conversations, we see AI providing an explicit answer to the question. We also find that the user is clearly adapting their answers based on the AI: the rate that user-AI is right is much higher when AI is right, than when AI is wrong or when the user is answering alone. Finally, we also see interesting differences across subjects: Moral Scenarios is distinguished by greater user independence and less reliance on AI, resulting in fewer near-exact rephrasings and explicit AI answers, and greater divergence between the AI answer being right and the user-AI answer being right.
@@
Statistic Math Conceptual Physics Moral Scenarios
Avg # user messages 1.57 1.37 1.69
Avg character length of user message 88.21 72.93 105.24
Avg # AI messages 1.48 1.32 1.51
Avg character length of AI message 793.92 918.40 1073.03
Conversation length: 2+ messages (%) 98% 98% 97%
Conversation length: 4+ messages (%) 27% 21% 37%
Conversation length: 6+ messages (%) 11% 8% 11%
First prompt: near-exact rephrase (%) 58% 62% 23%
First prompt: component (%) 25% 29% 63%
Explicit AI answer in conversation (%) 86% 82% 50%
AI answer right (%) 76% 84% 78%
AI answer right, user-AI right (%) 91% 91% 82%
AI answer wrong, user-AI right (%) 42% 32% 56%
Table B1: Statistics of user-AI conversations in ChatBench, computed over the full study.